# DiCoFlex: Model-agnostic diverse counterfactuals with flexible control

### v1

@@ A.1 Lending Club
The Lending Club dataset [12] contains detailed information about loans issued through the Lending Club peer-to-peer lending platform. It includes borrower characteristics (such as credit score, annual income, employment length), loan specifics (loan amount, interest rate, purpose), and performance indicators (payment status, delinquency). The binary classification task is to predict whether a loan will be fully paid or charged off (default). This dataset is particularly relevant for financial counterfactual explanations as it represents real-world credit risk assessment scenarios where understanding model decisions is crucial for both borrowers and lenders.
A.2 Give Me Some Credit
@@
The Give Me Some Credit dataset [25] contains anonymized records of credit users with features such as debt-to-income ratio, number of times delinquent, monthly income, age, and number of open credit lines. The target variable indicates whether a user experienced a serious delinquency (more than 90 days overdue) within the previous two years. This dataset provides insight into credit risk prediction in consumer finance, where counterfactual explanations can offer actionable guidance to consumers looking to improve their creditworthiness.
A.3 Bank Marketing
@@
The Bank Marketing dataset [18] contains information from a direct marketing campaign conducted by a Portuguese banking institution. The features include client data (age, job, marital status, education), campaign contact information (communication type, day, month), economic indicators, and previous campaign outcomes. The prediction task is to determine whether a client will subscribe to a term deposit. This dataset represents a real-world marketing scenario where understanding model decisions can improve campaign efficiency and provide insights for personalized marketing strategies.
A.4 Credit Default
@@
The Credit Default dataset [41] contains information on credit card clients in Taiwan, including demographic factors, credit data, payment history, and bill statements. The target variable indicates whether the client defaulted on their payment in the following month. With 23 features (14 numerical and 9 categorical), this dataset presents complex feature interdependencies common in financial data. The dataset is valuable for counterfactual explanation research because it represents real-world credit risk assessment with diverse feature types and nonlinear relationships.
A.5 Adult
@@
The Adult Census dataset [2] contains demographic information extracted from the 1994 U.S. Census database. Features include age, education, occupation, work hours per week, and capital gain/loss. The binary classification task is to predict whether an individual’s income exceeds \$50,000 per year. This dataset is widely used in fairness and explainability research, as it contains sensitive attributes like race, gender, and age, making it valuable for studying how counterfactual explanations handle demographic factors.
Appendix B Evaluation Metrics
@@
[…]
@@
Proximity $\text{Proximity Num}\ (\downarrow)=\frac{1}{N_{\text{val}}}\sum_{n=1}^{N_{\text{val}}}\|\mathbf{x}^{0}_{n,\text{num}}-\mathbf{x}^{\prime 1}_{n,\text{num}}\|_{1}$ (13)
@@
Diversity is measured by Hypervol. log scale [21], which quantifies spread and coverage in objective space.
@@
Appendix C Generative Model Selection
@@
We compared three normalizing flow architectures—MAF [23], NICE [6], and Real NVP [7], as well as KDE as a non-parametric baseline. MAF significantly outperformed all alternatives, while NICE and RealNVP showed comparable performance, and KDE exhibited the poorest results.
@@
Appendix D Details of training algorithm
@@
Appendix E Limitations of our method
@@
Appendix F Computational Resources
@@
Our experimental framework utilized Python [38] as the primary programming language, Additionally, the open-source machine learning library PyTorch [24] is used to implement DiCoFlex. All experiments were conducted on a GPU cluster equipped with a GeForce RTX 4090 graphics card (24 GB VRAM) and an AMD Ryzen Threadripper PRO 5975WX 32-core processor, with 256 GB of available RAM. These resources provide sufficient computational power and processing speed to meet the requirements of our algorithm.
Appendix G Additional experimental results
G.1 Runtime Comparison Across Datasets
@@
Figure 3 displays runtime comparisons between DiCoFlex and baseline methods across all datasets, with time shown on a logarithmic scale. Both variants of DiCoFlex consistently achieve substantially faster execution times for counterfactual generation, while competing approaches demand processing times that are orders of magnitude longer. This substantial performance advantage stems from fundamental architectural differences. DiCE [19] performs separate optimization procedures for each explanation. CCHVAE [25] requires expensive latent space searching. Similarly, Wachter [40], ReViSE [13] and TABCF [22] rely on iterative or gradient-based optimization procedures that scale poorly with the number of counterfactuals generated. In contrast, DiCoFlex leverages conditional normalizing flows trained solely on labeled data to generate multiple diverse counterfactuals in a single forward pass. By eliminating iterative optimization procedures and model access at inference time, DiCoFlex enables real-time counterfactual generation.
G.2 Impact of Actionability Constraints
@@
Table 5 presents the evaluation of counterfactual explanations generated by DiCoFlex with different actionability constraints imposed through feature masks. The analysis reveals complex relationships between masking constraints and evaluation metrics. In particular, the application of different masks results in varying impacts across metrics without following a consistent directional pattern.

### v2

@@ B.1 Lending Club
The Lending Club dataset [jagtiani2019roles] contains detailed information about loans issued through the Lending Club peer-to-peer lending platform. It includes borrower characteristics (such as credit score, annual income, employment length), loan specifics (loan amount, interest rate, purpose), and performance indicators (payment status, delinquency). The binary classification task is to predict whether a loan will be fully paid or charged off (default). This dataset is particularly relevant for financial counterfactual explanations as it represents real-world credit risk assessment scenarios where understanding model decisions is crucial for both borrowers and lenders.
B.2 Give Me Some Credit
@@
The Give Me Some Credit dataset [pawelczyk2020learning] contains anonymized records of credit users with features such as debt-to-income ratio, number of times delinquent, monthly income, age, and number of open credit lines. The target variable indicates whether a user experienced a serious delinquency (more than 90 days overdue) within the previous two years. This dataset provides insight into credit risk prediction in consumer finance, where counterfactual explanations can offer actionable guidance to consumers looking to improve their creditworthiness.
B.3 Bank Marketing
@@
The Bank Marketing dataset [moro2014data] contains information from a direct marketing campaign conducted by a Portuguese banking institution. The features include client data (age, job, marital status, education), campaign contact information (communication type, day, month), economic indicators, and previous campaign outcomes. The prediction task is to determine whether a client will subscribe to a term deposit. This dataset represents a real-world marketing scenario where understanding model decisions can improve campaign efficiency and provide insights for personalized marketing strategies.
B.4 Credit Default
@@
The Credit Default dataset [default_of_credit_card_clients_350] contains information on credit card clients in Taiwan, including demographic factors, credit data, payment history, and bill statements. The target variable indicates whether the client defaulted on their payment in the following month. With 23 features (14 numerical and 9 categorical), this dataset presents complex feature interdependencies common in financial data. The dataset is valuable for counterfactual explanation research because it represents real-world credit risk assessment with diverse feature types and nonlinear relationships.
B.5 Adult
@@
The Adult Census dataset [adult] contains demographic information extracted from the 1994 U.S. Census database. Features include age, education, occupation, work hours per week, and capital gain/loss. The binary classification task is to predict whether an individual’s income exceeds \$50,000 per year. This dataset is widely used in fairness and explainability research, as it contains sensitive attributes like race, gender, and age, making it valuable for studying how counterfactual explanations handle demographic factors.
Appendix C Evaluation Metrics
@@
$\text{Sparsity Cat/Num}\ (\downarrow)=\frac{1}{N_{\text{val}}}\sum_{n=1}^{N_{\text{val}}}\frac{\|\mathbf{x}^{0}_{n,\text{cat/num}}-\mathbf{x}^{\prime 1}_{n,\text{cat/num}}\|_{0}}{D_{\text{cat/num}}},$ (29)
@@
$\text{Proximity Num}\ (\downarrow)=\frac{1}{N_{\text{val}}}\sum_{n=1}^{N_{\text{val}}}\|\mathbf{x}^{0}_{n,\text{num}}-\mathbf{x}^{\prime 1}_{n,\text{num}}\|_{1}$ (31)
@@
Diversity is measured by Hypervol. log scale [nowak2014empirical], which quantifies spread and coverage in objective space.
@@
Appendix D Generative Model Selection
@@
We compared three normalizing flow architectures—MAF [PapamakariosMP17], NICE [DinhKB14], and Real NVP [DinhSB17], as well as KDE as a non-parametric baseline. MAF significantly outperformed all alternatives, while NICE and RealNVP showed comparable performance, and KDE exhibited the poorest results.
@@
Appendix E Details of training algorithm
@@
E.1 Training Details and Hyperparameter Exploration
We report additional details of the DiCoFlex training process and hyperparameter configuration used in all experiments. All experiments were conducted on a GPU workstation equipped with an NVIDIA RTX 4090 (24 GB VRAM) and an AMD Ryzen Threadripper PRO 5975WX CPU with 256 GB RAM. Each model was trained for a maximum of 1000 epochs using the Adam optimizer ($\text{lr}=10^{-4}$) and early stopping with a patience of 300 epochs based on the validation objective.
Training times.
Training times for all benchmark datasets are visualized in Figure 3. Each subfigure corresponds to one dataset (Adult, Bank Marketing, Default, Give Me Some Credit, and Lending Club) and reports the wall-clock time required to complete full training, including early stopping. The training time primarily depends on the dataset size and feature dimensionality.
Hyperparameter exploration.
We explored several key hyperparameters influencing model behavior and training stability:
  • Number of nearest neighbors $K\in\{8,\,16,\,32\}$
  • Actionability penalty $\alpha\in\{1,\,10,\,1000\}$
  • MAF hidden features $\in\{16,\,32,\,64\}$
  • MAF hidden layers $\in\{2,\,5\}$
For each combination, models were trained under identical conditions with early stopping. Hyperparameter selection was guided by validation performance averaged across validity, proximity, and sparsity metrics, favoring configurations that balanced these criteria without overfitting. Unless otherwise stated, the final configuration used K=16K{=}16, $\alpha{=}10$, 32 hidden features, and 5 hidden layers.
Sensitivity analysis.
To verify robustness, we performed a limited sensitivity analysis on the Adult dataset, varying each hyperparameter within the ranges listed above. Results confirmed that DiCoFlex remains stable across a wide range of $K$, $\alpha$, and model capacities, with minimal variation in validity and plausibility metrics. Detailed quantitative sensitivity results are provided in Appendix E.2.
Appendix F Limitations of our method
@@
Recent work has demonstrated that counterfactual explanations can be exploited for model extraction attacks, as they reveal decision boundary information [aivodji2020model, wang2022dualcf]. Furthermore, diverse counterfactual explanations can particularly enhance extraction effectiveness. However, DiCoFlex’s design may offer partial protection. Unlike methods exploring arbitrary feature space regions, the learned normalizing flow captures only the conditional distribution over plausible, proximal counterfactuals, potentially limiting information available for extraction. Nevertheless, some extraction risk remains. DiCoFlex’s diversity may still allow adversaries to triangulate decision boundaries from multiple sparse directions.
DiCoFlex generates counterfactuals based on proximity and plausibility without explicit fairness constraints. If the underlying model exhibits demographic biases, DiCoFlex may generate counterfactuals that reflect or amplify these biases. Future work should investigate incorporating fairness-aware constraints into the normalizing flow training to ensure protected attributes are handled appropriately.
As DiCoFlex trains on $K$-nearest neighbors from labeled training data (Equation 6), counterfactual quality depends on label accuracy.
Appendix G Context Vector Architecture
The conditional normalizing flow $p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime},p,\mathbf{m})$ requires a mechanism to incorporate conditioning information during the generative process. We achieve this through a context vector that concatenates all conditioning inputs:
$\mathbf{c}=[\mathbf{x},y^{\prime},p,\mathbf{m}]\in\mathbb{R}^{d+1+1+d},$ (32)
where $\mathbf{x}\in\mathbb{R}^{d}$ is the original instance, $y^{\prime}\in\{0,1\}$ is the target class (one-hot encoded for multi-class scenarios), $p\in\mathbb{R}$ is the sparsity parameter controlling the $L_{p}$ norm, and $\mathbf{m}\in\{0,1\}^{d}$ is the binary actionability mask indicating which features should remain unchanged.
Following the Masked Autoregressive Flow (MAF) architecture [PapamakariosMP17], this context vector $\mathbf{c}$ is passed to each layer of the flow through the affine coupling transformations. Specifically, in MAF, each transformation layer computes:
$\mathbf{z}_{i}=\mathbf{x}^{\prime}_{i}\odot\exp(\alpha_{i}(\mathbf{x}^{\prime}_{<i},\mathbf{c}))+\mu_{i}(\mathbf{x}^{\prime}_{<i},\mathbf{c}),$ (33)
where $\alpha_{i}$ and $\mu_{i}$ are neural networks that produce scale and shift parameters conditioned on both previous dimensions $\mathbf{x}^{\prime}_{<i}$ and the context vector $\mathbf{c}$. This autoregressive conditioning allows the flow to learn expressive transformations that respect the user-specified constraints encoded in $\mathbf{c}$.
The key advantage of this architecture is that it enables flexible control at inference time: users can adjust $p$ and $\mathbf{m}$ dynamically to generate counterfactuals satisfying different sparsity and actionability requirements without retraining the model. The flow learns to map these constraint specifications to appropriate regions of the counterfactual distribution during training, as described in Algorithm 1.
Appendix H Computational Resources
@@
Our experimental framework utilized Python [van1995python] as the primary programming language, Additionally, the open-source machine learning library PyTorch [pytorch] is used to implement DiCoFlex. All experiments were conducted on a GPU cluster equipped with a GeForce RTX 4090 graphics card (24 GB VRAM) and an AMD Ryzen Threadripper PRO 5975WX 32-core processor, with 256 GB of available RAM. These resources provide sufficient computational power and processing speed to meet the requirements of our algorithm.
Appendix I Additional experimental results
I.1 Runtime Comparison Across Datasets
@@
Figure 3 displays runtime comparisons between DiCoFlex and baseline methods across all datasets, with time shown on a logarithmic scale. Both variants of DiCoFlex consistently achieve substantially faster execution times for counterfactual generation, while competing approaches demand processing times that are orders of magnitude longer. This substantial performance advantage stems from fundamental architectural differences. DiCE [mothilal2020explaining] performs separate optimization procedures for each explanation. CCHVAE [pawelczyk2020learning] requires expensive latent space searching. Similarly, Wachter [Wachter2017counterfactual], ReViSE [joshi2019towards] and TABCF [panagiotou2024TABCF] rely on iterative or gradient-based optimization procedures that scale poorly with the number of counterfactuals generated. In contrast, DiCoFlex leverages conditional normalizing flows trained solely on labeled data to generate multiple diverse counterfactuals in a single forward pass. By eliminating iterative optimization procedures and model access at inference time, DiCoFlex enables real-time counterfactual generation.
I.2 Impact of Actionability Constraints
@@
Table˜6 presents the evaluation of counterfactual explanations generated by DiCoFlex with different actionability constraints imposed through feature masks. The analysis reveals complex relationships between masking constraints and evaluation metrics. In particular, the application of different masks results in varying impacts across metrics without following a consistent directional pattern.
@@
I.3 Statistical Uncertainty of Results
To assess the robustness of the reported metrics, we compute standard deviations across counterfactual sets generated per factual instance. For each test point, we produce multiple counterfactuals and evaluate the per-instance metric variance; the reported values represent the mean standard deviation across the test set. Note that the hypervolume metric operates at the set level and thus does not admit a standard deviation estimate.
Table 7: Standard deviations of evaluation metrics across datasets and methods. Values indicate variability across generated counterfactuals for each instance. Lower values correspond to more stable generation behavior. Standard deviations are shown for all metrics except Hypervolume, which is computed at the set level and thus not associated with instance-wise variance.
Validity ↑ Classif. Proximity Sparsity $\epsilon$-sparsity LOF
Dataset Model prob. ↑ cont. ↓ cat. ↓ cont. ↓ log scale ↓
Lending DiCoFlex (p=2.0) 0 0.012 0.439 0.198 0.087 0.076
Club DiCoFlex (p=0.01) 0 0.031 0.432 0.209 0.100 0.053
CCHVAE 0 0.368 0.401 0.049 0.066 0.119
DiCE 0 0.037 1.358 0.004 0.053 0.167
Adult DiCoFlex (p=2.0) 0 0.011 0.518 0.099 0.127 0.256
DiCoFlex (p=0.01) 0 0.000 0.352 0.106 0.096 0.812
CCHVAE 0 0.102 0.646 0.049 0.189 0.976
DiCE 0 0.000 1.464 0.017 0.264 0.915
Bank DiCoFlex (p=2.0) 0 0.025 0.571 0.100 0.111 0.016
DiCoFlex (p=0.01) 0 0.119 0.466 0.088 0.073 0.011
CCHVAE 0 0.053 0.606 0.061 0.090 0.078
DiCE 0 0.012 1.162 0.056 0.082 0.139
Default DiCoFlex (p=2.0) 0 0.021 0.346 0.138 0.071 0.038
DiCoFlex (p=0.01) 0 0.012 0.324 0.171 0.085 0.034
CCHVAE 0 0.043 0.273 0.097 0.076 0.047
DiCE 0 0.097 1.203 0.141 0.113 0.073
GMC DiCoFlex (p=2.0) 0 0.032 0.535 0.235 0.058 0.061
DiCoFlex (p=0.01) 0 0.037 0.694 0.229 0.059 0.099
CCHVAE 0 0.052 0.507 0.148 0.076 0.064
DiCE 0 0.072 1.021 0.048 0.073 0.019
Notation: Standard deviations are denoted directly as metric values in this table (no $\pm$ notation used for compactness). Metrics are averaged across counterfactuals for each instance and then aggregated across the test set.
I.4 German Credit Dataset Results
To evaluate the generalization of DiCoFlex to smaller tabular datasets, we conduct an additional experiment on the German Credit dataset (1,000 samples, 20 features, binary target). Despite the limited data availability, DiCoFlex maintains consistent performance across all evaluation metrics, demonstrating robustness in low-data regimes.
Table 8: Comparison of counterfactual generation methods on the German Credit dataset. Lower proximity, sparsity, $\epsilon$-sparsity, and LOF indicate better results; higher validity, classification probability, and hypervolume indicate improved performance.
Classif. Proximity Sparsity $\epsilon$-sparsity LOF Hypervol.
Dataset Model Validity ↑ prob. ↑ cont. ↓ cat. ↓ cont. ↓ log ↓ log ↑
German DiCoFlex (p=2.0) 1.000 0.833 0.639 0.553 0.756 0.0465 -6.030
Credit DiCoFlex (p=0.01) 1.000 0.833 0.647 0.556 0.750 0.0443 -6.115
CCHVAE 1.000 0.650 0.610 0.538 0.825 0.0747 -8.878
DiCE 0.954 0.808 1.143 0.257 0.756 0.1936 -7.456
ReViSE 0.673 0.731 0.841 0.523 0.742 0.1731
TABCF 1.000 0.912 0.829 0.348 0.559 0.0508
Wachter 0.356 0.508 1.633 0.000 0.952 0.2529
Even under constrained data conditions, both DiCoFlex variants (DiCoFlex (p=2.0) and DiCoFlex (p=0.01)) achieve perfect validity and strong plausibility (lowest LOF), confirming their ability to learn consistent counterfactual manifolds from limited examples. The results also show that the hypervolume metric remains competitive, suggesting that diversity is preserved even when the training data distribution is relatively sparse.
These findings highlight that DiCoFlex generalizes effectively to smaller datasets, further reinforcing its practicality for real-world applications where large annotated samples may be unavailable.